OpenStudy (anonymous):

solve the following equations: 2^(-100x)=(0.5)^(X-4) and 4^x * (1/2)^(3-2x)=8 * (2^x)^2

OpenStudy (solomonzelman):

\[2^{(-100x)}=(0.5)^{(X-4)}\] do you know how to use logarithms on this?

OpenStudy (anonymous):

i know to get an equivalent base and that is what i don't know how to get.

OpenStudy (solomonzelman):

\[a^B=c^F~~~~~~~->~~~~~~\log(a^B)=Log(c^F)\]do this process and use the\[\log(a^b)=b~\log~a\]

OpenStudy (anonymous):

is there a way without logs, that is a future lesson. not for this one.

OpenStudy (ranga):

yes. write 0.5 as 1/2 = 2^(-1) The bases will be the same on both sides and therefore the exponents will be the same.

OpenStudy (solomonzelman):

Man!! Haven't though of that!

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.